OpenStudy (anonymous):

find the solution of the equation in the interval notation[0,2pi) cos^3= cos x

OpenStudy (campbell_st):

well rewrite the equation as \[\cos^3(x) - \cos(x) = 0\] factoring you get \[\cos(x)(\cos^2(x) - 1) = 0\] factor the quadratic and you should be then able to identify all the solutions..

OpenStudy (anonymous):

can u pls check my answer, is x=pi,x=pi/2,x=3pi/2

OpenStudy (campbell_st):

well you should have found cos(x) = -1, 0 and 1 so the inverse are 0, pi/2, pi, 3pi/2

OpenStudy (anonymous):

your awesome. thank you.